OpenStudy (anonymous):

how do you solve. these of fractions... 5/n-3 + n/n+3

OpenStudy (anonymous):

Combine them and their denominator (5+n)/(n-3) Done lol

OpenStudy (anonymous):

really lol

hero (hero):

lol....no

hero (hero):

It's not that simple.

OpenStudy (anonymous):

yea i seen it was wrong lol i think i may have got it tho. just a lot of steps but after working about 5 i think i get it

hero (hero):

He doesn't even explain what happened to the n + 3

OpenStudy (anonymous):

yea i know

hero (hero):

\[\frac{5}{n - 3} + \frac{3}{n + 3}\]

hero (hero):

To add fractions like these, let's go back to actual numbers. Suppose you had to add \[\frac{2}{3} + \frac{4}{5}\] What would you do?

OpenStudy (anonymous):

find the lcd

OpenStudy (anonymous):

Oh sorry i didn't see it was n-3 and n+3 i thought they were both n-3 sorry

OpenStudy (anonymous):

This chapter was annoying man are you on flvs?

hero (hero):

You would realize that you need to multiply the first fraction by 5/5 and the second fraction by 3/3 \[\frac{2}{3} \times \frac{5}{5} + \frac{4}{5} \times \frac{3}{3}\] We do this because in order to add fractions we need both denominators to be the same.

OpenStudy (anonymous):

\[[(n+3)/(n+3)]5/(n-3) + n/(n+3)[(n-3)/(n-3)]\]

hero (hero):

Upon doing so we get \[\frac{10}{15} + \frac{12}{15}\] Which equals \[\frac{22}{15}\]

hero (hero):

The same process can be applied to the fractions you have to add.

OpenStudy (anonymous):

\[\frac{ 5n+15+n^2-3n }{ (n+3)(n-3) }\]

hero (hero):

\[\frac{5}{n - 3} \times \frac{n + 3}{n + 3} + \frac{n}{n + 3} \times \frac{n - 3}{n - 3}\]

OpenStudy (anonymous):

\[\frac{ n^2+2n+15 }{ (n+3)(n-3) }\]

OpenStudy (anonymous):

\[\frac{ (n+5)(n-3) }{ (n+3)(n-3) }\]

OpenStudy (anonymous):

cancel out the n-3's

OpenStudy (anonymous):

\[\frac{ n+5 }{ n+3 }\]

OpenStudy (anonymous):

that's the answer

OpenStudy (anonymous):

Get ITTTT ?

hero (hero):

Only thing is, it's still an improper fraction

OpenStudy (anonymous):

I know but in the textbook, anywhere that will be the answer

OpenStudy (anonymous):

#adding and subtracting rational expressions chapter 7 section 3 XDD

hero (hero):

I was going to finish by saying that \[\frac{22}{15}\] is an improper fraction. And you express the improper fraction in simpliest from as a mixed fraction: \[\frac{22}{15} = \frac{15 + 7}{15} = \frac{15}{15} + \frac{7}{15} = 1 \frac{7}{15}\]

hero (hero):

\[\frac{n + 5}{n + 3}\] is also an improper fraction. To express it as a mixed fraction, do the following: \[\frac{n + 5}{n + 3} = \frac{n + 3 + 2}{n + 3} = \frac{n + 3}{n + 3} + \frac{2}{n + 3} = 1 + \frac{2}{n + 3}\]